Circular No. 5900

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN OPHIUCHUS B. A. Harmon and G. J. Fishman, Marshall Space Flight Center,

NASA; W. S. Paciesas, University of Alabama, Huntsville; and S. N. Zhang, Universities Space Research Association, report for the Compton Observatory BATSE Team: “The hard x-ray transient GRO J1719-24 has been remarkable in its steady, persistent, high level of emission. It reached a peak intensity of ∼ 1.4 Crab (20-100 keV) five days after discovery (IAUC 5874) on Sept. 30. The 1/e decay time is ∼ 300 days from a fit of occultation data obtained during Oct. 1-Nov. 22, although a linear fit to the light curve is equally good, with intensity dropping at ∼ 0.3 ± 0.05 percent per day relative to the peak. This is believed to be the slowest decay of any hard x-ray nova observed. Power-law fits to the spectrum in the band 20-100 keV showed the spectrum softening from −2.0 to −2.3 (± 0.05) during the rise, then softening more gradually since the end of September; it continues to be the brightest source in the sky in this energy range.”

GRS 1915+105 L. F. Rodriguez, National Radio Astronomy Observatory and Insti-

tuto de Astronomía, Universidad Nacional Autónoma de México; and I. F. Mirabel, Service d’Astrophysique, Centre d’Etudes de Saclay, report: “Using the Very Large Array, we are observing a strong burst in the radio counterpart of the hard x-ray transient in Aquila; on Nov. 29, we detected GRS 1915+105 in a remarkably high radio state. In our previous VLA observations of this source, it usually exhibited time-variable centimeter continuum emission at levels of a few mJy (IAUC 5773, 5830). Recent flux densities: Nov. 29, 90 mJy at 6 cm, 120 mJy at 3.5 cm, 130 mJy at 2 cm; Dec. 5, 300 mJy at 20 cm, 590 mJy at 6 cm, 450 mJy at 3.5 cm, 310 mJy at 2 cm, 190 mJy at 1.3 cm. The inverted spectrum observed on Nov. 29 is suggestive of optically-thick synchrotron emission, while that of Dec. 5 was indicative of optically-thin synchrotron emission for wavelengths below 6 cm and with a possible self-absorption turnover between 20 and 6 cm. It is of interest to obtain information at other wavelengths during this radio outburst. The position of the radio source is α = 19^{h}12^{m}49^{s}.97, δ = +10^{o}51′26′′.8 (equinox 1950.0; accuracy ∼ 1′′).”

1993 December 6 Daniel W. E. Green